OpenStudy (anonymous):

(-3) to the power of 5 please show steps

OpenStudy (anonymous):

-3^5= -3*-3*-3*-3*-3

OpenStudy (saifoo.khan):

-3 * -3 *-3 * -3 * -3

OpenStudy (aroub):

(-3)^5 -3*-3*-3*-3*-3=-243

OpenStudy (anonymous):

-3*-3*-3*-3*-3= 9*9*-3

OpenStudy (anonymous):

9*9*-3= 81*-3

OpenStudy (anonymous):

81*-3=-243